diffrentiate wiyh respect to x


(xcosx)^1/x+(xsinx)^x

Grade:

1 Answers

Chetan Mandayam Nayakar
11 years ago

let y=(xcosx)^1/x,ln(y)=(1/x)(ln(xcosx)),differentiate both sides w.r.t. x

let z=(xsinx)^x, ln(z)=xln(xsinx),differentiate both sides w.r.t. x

contact me if you need more explanation